OpenStudy (bekkah323):

x+y/xy^2 plus 3x+y/x^2y *equation below

OpenStudy (bekkah323):

\[\frac{ x+y }{ xy^2 }+\frac{ 3x+y }{ x^2y}\]

OpenStudy (anonymous):

to simplify this, start by putting both fractions under the same denominator which is (xy)^2

OpenStudy (bekkah323):

\[\frac{ x^2(x+y) }{ x^2y^2 }+\frac{ y^2(3x+y) }{ x^2y^2 }\]

OpenStudy (anonymous):

There's only one 'x' and 'y' outside the fraction as you're multiplying both numerator and denominator by 'x'/'y' \[\frac{ x(x+y) }{ x^2y^2 }+\frac{ y(3x+y) }{x^2y^2 }\]

OpenStudy (anonymous):

Then combine both fractions under one common denominator :)

OpenStudy (bekkah323):

oh, ok so would it be\[\frac{ x^2+4xy+y^2 }{ x^2y^2 }\]

OpenStudy (souvik):

right!